2. An automobile tyre of volume 28dm3

is filled with air at a pressure of 1 atm.What volume does this air

occupy at higher pressure 2.5 atm with the same temperature?
1
Expert's answer
2020-06-09T07:18:59-0400

Solution.

"p1 \times V1 = p2 \times V2"

"V2 = \frac{p1 \times V1}{p2}"

V2 = 11.2 dm3

Answer:

V2 = 11.2 dm3